# Substrate Scope Contrastive Learning: Repurposing Human Bias to Learn Atomic Representations

### v1

@@ 10 Details of reactivity descriptors
Figure S5: All computed global (molecule-level) descriptors using autoqchem 57, and an explanation of each.
Figure S6: All computed local (atom-level) descriptors using autoqchem57, and an explanation of each.
@@ 11 Additional Results
11.1 Intuitive investigation of the learned embeddings
To offer an intuitive examination of the bit-level details of the embeddings, we illustrate the bit value of the corresponding position in the embeddings with a heatmap and their hierarchical clustering in Figure S7. We analyzed the same set of aryl bromides as Figure 2B-C.
Figure S7: Hierarchical clustering analysis of learned embeddings for a set of aryl bromide molecules. The dendrogram on the topindicates the similarity between the aryl bromide variants. The heatmap below shows the bit value of the corresponding position in the embeddings for each molecule.
@@ 11.2 Visualization of learned aryl halide chemical spaces
To offer insights into the evolution of the embeddings during training, we present the comparative analysis of our embeddings using PCA, t-SNE, and UMAP projections, both pre- and post-training, in Figure S8. This analysis utilizes the identical set of molecules featured in Figure 2A.
@@ 11.3 Correlation with conventional reactivity indicators
Figure S12: The t-SNE projection colored with various reactivity descriptors.
Figure S13: The t-SNE projection colored with various reactivity descriptors. (Continued)
@@ 11.4 Regression performance on the training substrate scopes
We tried to use the learned embedding as feature to predict reaction yields in the training data. Each substrate scope was treated as a single regression task and leave-one-out validation $r^{2}$ for each scope is shown. The methods compared include ContraScope combined with k-Nearest Neighbors (kNN), RDKit2D with kNN, and RDKit2D with Random Forest (RF). The distributions highlight the variability and challenges encountered in the predictive modeling of chemical yields, with all approaches showing a wide distribution of $r^{2}$ values, including negative values indicative of a failure to capture any trend.
Figure S14: Distribution of the $r^{2}$ values for yield prediction across training substrate scopes using nearest neighbor regression.
@@ 11.5 Investigation on chemistry informers
In the study by Kutchukian et al. (2016)61, a library of 18 aryl halides was utilized to explore their reactivity in the Buchwald-Hartwig reaction under 18 distinct conditions. We encoded these aryl halides from the specified aryl halide informer library. Subsequently, our analysis involved a comparative assessment of the pair-wise signal-to-noise ratio (SNR) distances derived from the ContraScope embeddings against the pair-wise Euclidean distances computed from the vectors of reported yields. This comparison is graphically represented using heatmaps in Figure S15. However, it is noteworthy that this analysis did not reveal congruent patterns in the left and right panels of the figure.
@@
11.6 Less effective in learning global reactivity contribution
Similar to Figure 3B, we evaluated the regression performance for various molecular embeddings in predicting global chemical reactivity descriptors. A comparison between the performance in global and local descriptors are shown in Figure S16. The results show that while common embeddings yield moderate $r^{2}$ values across the descriptors, our embedding exhibits a lower performance, suggesting a less effective capture of global reactivity information in our model.
Figure S16: A comparative study of evaluation of regression performance for various molecular embeddings in predicting global and local chemical reactivity descriptors.

### v2

@@ 11 Details of reactivity descriptors
Figure S12: All computed global (molecule-level) descriptors using auto-qchem 24 and an explanation of each.
Figure S13: All computed local (atom-level) descriptors using auto-qchem24 and an explanation of each.
@@ 12 Additional Results
12.1 Qualitative investigation of the learned embeddings
To offer an intuitive examination of the bit-level details of the embeddings, we illustrate the bit value of the corresponding position in the embeddings with a heatmap and their hierarchical clustering in Figure S14. We analyzed a set of 12 hand-selected para-substituted bromobenzenes, as shown in the figure.
In order to validate whether quantitative distances in the learned embedding space accurately reflect known functional similarity in chemical reactivity, we next encoded the same set of para-substituted bromobenzenes, then calculated and visualized their pairwise distances (Figure S15). The heatmap reveals that molecules with electron-withdrawing groups, such as nitrile, ester, and trifluoromethyl, demonstrate notably reduced distances compared to those with electron-donating groups like methoxy, hydroxyl, ether, and amine. This discernible pattern is not mirrored in the Tanimoto distances55, 56 (refer to Figure S15), where distances are measured based on shared structural features in Morgan fingerprints. This comparison emphasizes that distances in our learned embedding space more effectively capture functional similarities in chemical reactivity than standard measures of structural similarity.
Figure S14: Heatmap of the pair-wise distance between 12 para-substituted aryl bromides, showcasing that the learned embeddings’ signal-to-noise (SNR) distances (A) better reflects functional similarity than the conventional Tanimoto distances based on structural fingerprints (B).
Figure S15: Pairwise distance heatmaps for 12 para-substituted bromobenzenes. (A) Distances calculated using ContraScope embeddings reveal clear clustering of molecules based on functional reactivity, with electron-withdrawing groups showing reduced distances compared to electron-donating groups. (B) Distances derived from Morgan fingerprints using Tanimoto similarity do not exhibit this functional clustering, instead reflecting structural similarity.
@@ 12.2 Visualization of learned aryl halide chemical spaces
To offer insights into the evolution of the embeddings during training, we present the comparative analysis of our embeddings using PCA, t-SNE, and UMAP projections, both pre- and post-training, in Figure S16. This analysis utilizes the identical set of molecules featured in Figure 2.
@@ 12.3 Correlation with conventional reactivity indicators
Figure S20: The t-SNE projection of learned embeddings with each point colored with various reactivity descriptors.
Figure S21: The t-SNE projection of learned embeddings with each point colored with various reactivity descriptors. (Continued)
@@ 12.4 Regression performance on the training substrate scopes
We tried to use the learned embedding as feature to predict reaction yields in the training data. Each substrate scope was treated as a single regression task and leave-one-out cross validation $r^{2}$ for each scope is shown. The methods compared include ContraScope combined with k-Nearest Neighbors (kNN), RDKit2D with kNN, and RDKit2D with Random Forest (RF). The distributions highlight the variability and challenges encountered in the predictive modeling of chemical yields, with all approaches showing a wide distribution of $r^{2}$ values, including negative values indicative of a failure to capture any trend.
Figure S22: Distribution of the $r^{2}$ values for yield prediction across all substrate scopes using a leave-one-out cross-validation to evaluate each. None of the representations tested provides a promising degree of predictivity.
@@ 12.5 Investigation on chemistry informers
In the study by Kutchukian et al. (2016)32, a library of 18 aryl halides was utilized to explore their reactivity in the Buchwald-Hartwig reaction under 18 distinct conditions. We encoded these aryl halides from the specified aryl halide informer library. Subsequently, our analysis involved a comparative assessment of the pair-wise signal-to-noise ratio (SNR) distances derived from the ContraScope embeddings against the pair-wise Euclidean distances computed from the vectors of reported yields. This comparison is graphically represented using heatmaps in Figure S23. However, it is noteworthy that this analysis did not reveal congruent patterns in the left and right panels of the figure.
@@
12.6 Modified training strategies
To address potential biases in our negative sampling strategy, we explored hard negative sampling approaches. First, we focused on substrates with the same halogen type when generating negative samples, treating substrates excluded from the published scope as negative examples only if their halogen matched that of the substrates selected as the anchor. This adjustment aimed to mitigate differences in halogen reactivity, ensuring that the generated negative examples reflected comparable chemical contexts and reducing confounding effects. Additionally, we narrowed our focus to Pd-catalyzed reactions, based on the NameRXN reaction category, to eliminate noise arising from varying reactivities required for different types of reactions. Despite these efforts, neither strategy yielded meaningful improvements.
The impact of this strategy is illustrated in the learning curves shown in Figure S24. The results indicate that while the anchor-positive loss remains at a similar level and trend compared to the original sampling strategy, the anchor-negative loss is shifted upward. This shift suggests that the hard negative sampling strategy effectively introduces more challenging examples for the model to distinguish, validating its intended purpose. However, despite the effective negative sampling, we did not observe a quantitative improvement in performance metrics. We hypothesize that this arises from the limited modeling capacity of the specific network architecture employed in this study. This observation highlights the potential need for more expressive model architectures to fully leverage the benefits of hard negative sampling. We report the model trained with the original strategy in the main text and leave further exploration for future work.
(a) Total substrate scope contrastive loss.
(b) The value of the anchor-positive term.
(c) The value of the anchor-negative term.
Figure S24: The comparison of learning curves of the pre-training process with (Same halogen negative) and without (Previous loss) same halogen sampling.
@@ 12.7 Substrate scope comparison
Substrate scope selection. We illustrate how learned embeddings can assist in the selection of a “diverse” set of aryl bromides, which one may wish to do at the outset of an experimental screening campaign. This process, inspired by Kutchukian et al.’s32 and Kariofillis et al.’s8 methodology, involved clustering the chemical space of commercially-available aryl bromides using K-means57; while the authors used a DFT-derived feature vector, here we use our learned embeddings. A representative aryl bromide is chosen from each cluster to form a set designed to exhibit a diverse range of reactivities (Figure S25). Unlike clustering based on structural fingerprints or expert-selected descriptors, our approach relies on the specific reactivity profiles of aryl halides as exemplified by our embeddings, in principle leading to a more curated set for this class of compounds better aligned with their unique reactivity characteristics. As there is no objective correct answer for diverse substrate selection, we leave the selection for qualitative interpretation by the reader.